# ResiDual Transformer Alignment with Spectral Decomposition

### v1

@@ 3.3 Spectral Dataset Comparison
The original principal angles measure the alignment between subspaces by maximizing cosine similarity of vectors in their span. In our discrete case, vectors are selected from sets ${\mathbb{S}}_{1}$ and ${\mathbb{S}}_{2}$ directly, with optional weighting. The final measure is the aggregation of the spectral cosine similarities along the $\min(k_{1},k_{2})$ entries, with normalization to bound our measure between 0 and 1. We define the normalized spectral cosine similarity between the two sets of principal components as:
@@
We consider the same ViT-based encoders of Section 3.1, and 14 different datasets: ImageNet (the same split used in Section 3.1), CIFAR(-100/-10) (Krizhevsky, 2009), ImageNet-Sketch (Wang et al., 2019), Cars (Krause et al., 2013), MNIST (LeCun et al., 1998), SVHN (Netzer et al., 2011), EuroSAT (Helber et al., 2019), RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2016), GTSRB (Stallkamp et al., 2011), PACS (Li et al., 2017), and random images (10000 samples with RGB values in $[-1,1]$). We use the original train/validation/test splits if available, otherwise we produce the splits through a stratified random sampling over the classes. For each encoder, we use our similarity measure to compare its unit representations produced on each training dataset with the ones obtained on the training split of ImageNet. ImageNet is taken as a reference under the assumption that, being a general enough dataset, its head PCA bases are sufficiently comprehensive to approximate the primary features across other datasets. Additionally, we perform a qualitative inspection of a few heads that stand out by finding their textual decomposition. For this step, we use Simultaneous Orthogonal Matching Pursuit (SOMP), having established its strong relationship with TextSpan (Section A.2).
@@
The results of the spectral head-to-head comparison between ImageNet and all other datasets on OpenCLIP-L are reported in Figure 4 (results on other models can be found in the Appendix in fig. 9). Rows are ordered according to the mean overall similarity between the corresponding dataset and ImageNet. Interestingly, dataset ordering is consistent across different encoders. The mean correlation coefficient between dataset similarities (averaged over all heads) across different models is 0.97. The full comparison between encoders is reported in the Appendix (Figure 9). On OpenCLIP-L, we observe that datasets that maximally align with ImageNet share classes with it (e.g., SUN397 and Sketch) and/or contain generic images (e.g., DTD and Cars). Moreover, datasets that share the same input image structure and concepts (CIFAR-10 and CIFAR-100) have an almost identical similarity distribution across heads. Overall, we observe a decreasing trend in head similarity scores as depth in the model increases. The simple, linear heads of the first layers are responsible for the extraction of low-level patterns (Dosovitskiy et al., 2021) and emerge as almost always identical across different data distributions. On the last layers, just a few heads per dataset stand out: this is where we are looking for specialization. Zooming in on a few of these heads, in Figure 4(b) we report their textual descriptions obtained with SOMP. Head 7 of layer 22 (specialized on seasons) stands out because it is highly activated in many datasets that contain pictures of scenery (such as SUN397, GTSRB, and more prominently EuroSAT). Head 11 of layer 22 (specialized on shades of gray) emerges as extremely different between ImageNet and Sketch, that contains grayscale drawings of ImageNet classes. Head 10 of layer 23 (specialized on numbers) is highly activated on both MNIST and SVHN. We note that this is not the only ’shared’ head between the two, but others, like head 1 of the same layer, are also activated by the random dataset, signaling that they are not as specific.
@@
From these findings, we observe that the similarity measure yields intuitive scores, with ImageNet’s foundational attributes demonstrating generalizability across various data distributions.
@@ 4 ResiDual Alignment
With a refined understanding of head specialization, our objective is now to leverage this property to enhance alignment between visual unit representations (both heads and MLPs) and text encodings in CLIP-like models. From this point onwards, our experiments will have a shared objective: given a zero-shot text classifier, i.e., text encodings for classes, we manipulate the residual to improve its alignment with the text subspace. Improved alignment directly benefits multimodal tasks such as zero-shot classification, as it strengthens the model capacity to interpret visual data through text-based descriptors.
@@ 4.1 Coarse Unit Alignment
We start by exploring whether certain heads are already aligned with the text subspace of interest for our task. Specifically, we aim to optimize vision-text alignment by combining visual heads selected using various scoring functions.
@@
We have 3 different selection methods: i) Unsupervised (U): We use the head-to-output correlation as a measure. Intuitively, the more a head correlates to the full output, the more information it carries. In practice, we compute the Pearson correlation between each sample at the head level and its corresponding output encoding, averaging across samples to obtain a scalar; ii) Task-conditioned Unsupervised (U|T): Conditioning on the output alone does not necessarily imply that the selected heads will be suitable for a given task. Since the task is modeled by a text subspace, having one encoding for each class, we can condition the previous unsupervised measure to be applied only on the head and output subspaces spanned by the task encodings. This has the effect of ignoring features that might influence the correlation but are not related to the task at hand. This is a direct application of the CompAttribute metric introduced in Balasubramanian et al. (2024); iii) Supervised (S): When we assume the availability not only of the task encodings, but also of labeled samples, we can directly estimate the head score by looking at its performance on the downstream task (in the style of logit lens (nostalgebraist, 2020)).
For each scoring function, we evaluate each head individually, rank them according to their scores, and apply a greedy top-k selection. We then sum these selected heads ${\mathbb{H}}^{\prime}\subseteq{\mathbb{H}}$ to create a partially recovered residual, which is subsequently evaluated on downstream task performance.
${\bm{Y}}^{\prime}=\sum_{i=1}^{|{\mathbb{H}}^{\prime}|}{\bm{H}}_{i}$ (5)
We have 3 control measures in place: i) Heads (H): The performance of the model when all the attention heads, and only them, are used. This gives information about the heads’ contribution to the residual and, symmetrically, how much the final performance depend on the MLP units; ii) Random (R): The average performance over 10 independent random samplings of $k$ head units. This can be seen as a lower bound on the expected performance; iii) Base (B): The original performance of the model without any modification to its residual. Intuitively, this could represent a theoretical upper bound on the performance if there are no task-aligned units.
@@
We evaluate these unit selection strategies on two CLIP-like models (BLIP-L and OpenCLIP-L) and 10 of the datasets of Section 3.3. We choose $k$ for the greedy and random selection so that 5% of the total heads are considered. Additional results are presented in the Appendix in Table 3, Table 4 and Table 5.
@@
As reported in Table 1, the unsupervised scoring (U) performs unexpectedly well, despite not explicitly considering the task. This effectiveness likely stems from the fact that core task information is often embedded in the first few principal components (PCs) of the output. Underlying this method is the assumption that first principal components already align with the main task-relevant information. By aligning with this information using only a few heads, we achieve a dual benefit: preserving essential task-relevant information while effectively filtering out noise. In fact, adding the conditioning on the task (U|T) is just slightly beneficial in terms of performance, with the exception of SVHN. The vast majority (on average around 90%) of alignment between task and residual comes from the head contributions, as witnessed by the similarity between the columns H and B. The optimized selection strategy (O) is extremely powerful, having the best score among them, and sometimes almost doubling the original model performances (B) showing how task-relevant information is already present in the residual, just hidden.
@@ 4.2 Spectral ResiDual Alignment
Given that: i) retaining only task-aligned units is beneficial for image-text alignment (Section 4.1); ii) unit specialization is essentially encoded in their principal components (Section 3.2); iii) components of general enough data distributions (e.g., ImageNet) capture specialized behaviour even on other datasets (Section 3.3), we propose a method to directly filter information along the residual at the spectral level: ResiDual.
In short, we start from the decomposition formula for the residual stream (Equation 1), and allow anisotropic scaling of each unit representation. Specifically, given a unit representation ${\bm{X}}$, its corresponding principal component basis $\bm{\Phi}$, and associated mean $\bm{\mu}$, we define the ResiDual transformation of ${\bm{X}}$ as:
$\text{RD}_{\bm{\Phi},\bm{\mu}}({\bm{X}},\bm{\lambda})=\mathbf{\Phi}^{T}\text{diag}(\bm{\lambda})\mathbf{\Phi}({\bm{X}}-\bm{\mu})^{T}$ (6)
@@
${\bm{Y}}^{\prime}=\sum_{i=1}^{|{\mathbb{U}}|}\text{RD}_{\bm{\Phi_{i}},\bm{\mu_{i}}}({\bm{U}}_{i},\bm{\lambda}_{i})$ (7)
In summary, ResiDual models a simple spectral anisotropic scaling of residual units, that results in more complex dynamics in the output space. In this section, we extensively evaluate the effectiveness of this method across a variety of configurations, models and datasets.
@@
We evaluate ResiDual in 3 different configurations: i) RD, as presented in the original formulation; ii) RD, a version of ResiDual constrained on the number and type of units (we restrict it to heads) and number of considered components (we truncate PCA bases to explain 90% of the variance). This is done because head units, and specifically their principal components, are strongly connected to specialization; iii) RDY, where the ResiDual transformation is applied directly on the output encoding ${\bm{Y}}$, to assess whether output components can already be aligned with the task. For all configurations, we select ImageNet as a reference for the PCA bases $\bm{\Phi}$ and unit means $\bm{\mu}$ that appear in the ResiDual Equation 6.
@@
For this experiment, we work with 3 CLIP-like models, BLIP-L, CLIP-L and OpenCLIP-L (CLIP/OpenCLIP-B results can be found in the Appendix in Table 6 and Figure 16), and tune them on the 10 datasets employed in Section 4.1. All training runs use the Schedule-Free Adam optimizer (Defazio et al., 2024) with the automatic learning rate finder by Smith (2017), implemented in PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). The maximum number of epochs is 30, with an early-stopping policy on the validation set accuracy with a patience of 5 epochs.
@@
Result analysis
A comparative analysis of ResiDual against reference alignment methods is reported in Figure 5. We observe that a linear transformation of the output is sufficient to approximate full finetuning performance. The spectral residual transformation modeled by ResiDual attains comparable (if not better) results than the linear aligner on all datasets. These statements hold true across all models. Specifically, the case of SVHN stands out: on this input dataset, ResiDual has an advantage of approximately 10% on all models. We hypothesize that this gap is due to the absence of task-relevant features at the output level. This is confirmed by the results in Table 2: on SVHN, RDY has the largest gap from RD, meaning that output components are not well aligned with the task. While having approximately 30% of the learnable parameters, RD achieves comparable results to RD, indicating that applying the ResiDual procedure to heads (and not MLPs) and their first principal components alone is enough. Moreover, we note that in the cases where the performance of the original model is already satisfactory (e.g., CIFAR-10), applying ResiDual does not compromise alignment.
@@ 5 Conclusions
ResiDual fundamentally works by extracting information from residual units already containing task-relevant principal components. We expect that when this assumption does not hold, ResiDual cannot recover the alignment, resulting in a significant drop in downstream performance. Moreover, our downstream tasks focused on zero-shot classification in CLIP-like models. This implies considering only the alignment between [CLS] tokens, ignoring sequence-level information.
@@
The ResiDual formulation is based solely on the residual decomposition technique, which opens to its application across virtually any transformer architecture. Our findings show that ResiDual can be limited to operating on a subset of residual units (i.e., attention heads). An additional constraint on selecting only contributions from the first few model layers creates opportunities to enhance model inference.
@@ A.1 Residual decomposition
Transformers are residual networks, made of stacked blocks that contain a Multi-Head Attention (MHA) layer and a MLP, both surrounded by a skip (residual connection). All vision transformers we employ in this work abide by the ’pre-norm’ Wang et al. (2019) architecture, which slightly modified the original (Vaswani et al., 2017) by moving LayerNorm before MHA and MLP sub-layers (Figure 6). This modification implies that, at each layer, MHA and MLP sub-layers directly write their output representation to the residual stream. Hence, the final latent representation of the model is given by:
@@
In many transformer models, such as the ones employed in this work, the final residual encoding ${\bm{Y}}$ is not the final output of the model. For instance, in CLIP, ${\bm{Y}}$ is passed through a LayerNorm and then through a biasless linear projection ${\bm{P}}$, that maps ViT encodings to the shared vision-language space:

### v2

@@ 3.3 Spectral Dataset Comparison
The original principal angles measure the alignment between subspaces by maximizing the cosine similarity of vectors in their span. In our discrete case, vectors are directly selected from sets ${\mathbb{S}}_{1}$ and ${\mathbb{S}}_{2}$ with optional weighting. The final measure is the aggregation of the spectral cosine similarities along the $\min(k_{1},k_{2})$ entries, with normalization to bound our measure between 0 and 1. We define the normalized spectral cosine similarity between the two sets of principal components as:
@@
We consider the same ViT-based encoders of Section 3.1, and 14 different datasets: ImageNet (the same split used in Section 3.1), CIFAR(-100/-10) (Krizhevsky, 2009), ImageNet-Sketch (Wang et al., 2019), Cars (Krause et al., 2013), MNIST (LeCun et al., 1998), SVHN (Netzer et al., 2011), EuroSAT (Helber et al., 2019), RESISC45 (Cheng et al., 2017), DTD (Cimpoi et al., 2014), SUN397 (Xiao et al., 2016), GTSRB (Stallkamp et al., 2011), PACS (Li et al., 2017), and random images (10000 samples with RGB values in $[-1,1]$). We use the original train/validation/test splits if available, otherwise we produce the splits through a stratified random sampling over the classes. For each encoder, we use our similarity measure to compare its unit representations produced on each training dataset with the ones obtained on the training split of ImageNet. ImageNet is taken as a reference under the assumption that being a general enough dataset, its head PCA bases are sufficiently comprehensive to approximate the primary features across other datasets. Additionally, we perform a qualitative inspection of a few heads that stand out by finding their textual decomposition. For this step, we use Simultaneous Orthogonal Matching Pursuit (SOMP), having established its strong relationship with TextSpan (Section A.2).
@@
The results of the spectral head-to-head comparison between ImageNet and all other datasets on OpenCLIP-L are reported in Figure 4 (results on other models can be found in the Appendix in fig. 13). Rows are ordered according to the mean overall similarity between the corresponding dataset and ImageNet. Interestingly, dataset ordering is consistent across different encoders. The mean correlation coefficient between dataset similarities (averaged over all heads) across different models is 0.97. The full comparison between encoders is reported in the Appendix (Figure 13). On OpenCLIP-L, we observe that datasets that maximally align with ImageNet share classes with it (e.g., SUN397 and Sketch) and/or contain generic images (e.g., DTD and Cars). Moreover, datasets that share the same input image structure and concepts (CIFAR-10 and CIFAR-100) have an almost identical similarity distribution across heads. Overall, we observe a decreasing trend in head similarity scores as depth in the model increases. The simple, linear heads of the first layers are responsible for the extraction of low-level patterns (Dosovitskiy et al., 2021) and emerge as almost always identical across different data distributions. On the last layers, just a few heads per dataset stand out: this is where we are looking for specialization. Zooming in on a few of these heads, in Figure 4(b), we report their textual descriptions obtained with SOMP. Head 7 of layer 22 (specialized on seasons) stands out because it is highly activated in many datasets that contain pictures of scenery (such as SUN397, GTSRB, and, more prominently, EuroSAT). Head 11 of layer 22 (specialized in shades of gray) emerges as extremely different between ImageNet and Sketch, which contains grayscale drawings of ImageNet classes. Head 10 of layer 23 (specialized in numbers) is highly activated on both MNIST and SVHN. We note that this is not the only ’shared’ head between the two, but others, like head 1 of the same layer, are also activated by the random dataset, signaling that they are not as specific.
@@
These findings show that the similarity measure yields intuitive scores, with ImageNet’s foundational attributes demonstrating generalizability across various data distributions.
@@ 4 ResiDual Alignment
With a refined understanding of head specialization, our objective is now to leverage this property to enhance alignment between visual unit representations (both heads and MLPs) and text encodings in CLIP-like models. From this point onwards, our experiments will have a shared objective: given a frozen zero-shot classifier, e.g., text encodings or prototypes for classes, we manipulate the residual to improve its alignment with this target subspace. In the main text, we focus on multimodal architectures, where improved alignment directly benefits tasks such as zero-shot classification by strengthening the model’s capacity to interpret visual data through text-based descriptors. Additionally, in Section A.3, we conduct experiments on unimodal encoders, where the zero-shot classifier is built in a prototype-based fashion.
@@ 4.1 Coarse Unit Alignment
We start by exploring whether certain heads are already aligned with the text subspace relevant to our task. Specifically, our goal is to optimize vision-text alignment by combining visual heads weighted using various scoring functions. These functions assign weights to entire head representations to modulate (in some cases completely remove) their contributions to the residual stream. Crucially, these methods operate exclusively at the level of whole-head reweighting, leaving the internal structure and intrinsic specialization of individual heads unchanged. As a result, only the output specialization is affected.
@@
We have 3 different selection methods: i) Unsupervised (U): We use the head-to-output correlation as a measure. Intuitively, the more one head correlates to the full output, the more information it carries. In practice, we compute the Pearson correlation between each sample at the head level and its corresponding output encoding, averaging across samples to obtain a scalar; ii) Task-conditioned Unsupervised (U|T): Conditioning on the output alone does not necessarily imply that the selected heads will be suitable for a given task. Since the task is modeled by a text subspace, having one encoding for each class, we can condition the previous unsupervised measure to be applied only on the head and output subspaces spanned by the task encodings. This has the effect of ignoring features that might influence the correlation but are not related to the task at hand. This is a direct application of the CompAttribute metric introduced in Balasubramanian et al. (2024); iii) Supervised (S): When we assume the availability not only of the task encodings but also of labeled samples, we can directly estimate the head score by looking at its performance on the downstream task (in the style of logit lens (nostalgebraist, 2020)).
For each scoring function, we evaluate each head individually, rank them according to their scores, and apply a greedy top-k selection. We then sum these selected heads ${\mathbb{H}}^{\prime}\subseteq{\mathbb{H}}$ to create a partially recovered residual, which is subsequently evaluated on downstream task performance:
${\bm{Y}}^{\prime}=\sum_{i=1}^{|{\mathbb{H}}^{\prime}|}{\bm{H}}_{i}\,.$ (5)
We have 3 control measures in place: i) Heads (H): The performance of the model when all the attention heads, and only them, are used. This gives information about the heads’ contribution to the residual and, symmetrically, how much the final performance depends on the MLP units; ii) Random (R): The average performance over 10 independent random samplings of $k$ head units. This can be seen as a lower bound on the expected performance; iii) Base (B): The original performance of the model without any modification to its residual. Intuitively, this could represent a theoretical upper bound on the performance if there are no task-aligned units.
@@
We evaluate these unit selection strategies on two CLIP-like models (BLIP-L and OpenCLIP-L) and 10 of the datasets of Section 3.3. We choose $k$ for the greedy and random selection so that 5% of the total heads are considered. Additional results are presented in the Appendix in Table 7, Table 4 and Table 5 for other CLIP-like models and Section A.3 for unimodal ones (DINOv2-L and ViT-L).
@@
Table 1: Accuracy when doing zero ablation of all units except top 5% of attention heads. Heads are assigned a binary weight using an Unsupervised (U), Task-conditioned Unsupervised (U|T), Supervised (S), and Random (R) strategy (a mean over 10 different seeds). H corresponds to using all the attention heads available, B is the original model performance, and O is the optimized continuous weighting case. For a qualitative analysis complementing these quantitative results on a selection of datasets, please refer to Table 6 in the Appendix.
@@
As reported in Table 1, the unsupervised scoring (U) performs unexpectedly well despite not explicitly considering the task. This effectiveness likely stems from the fact that core task information is often embedded in the first few principal components (PCs) of the output. By aligning with this information using only a few heads, we achieve a dual benefit: preserving essential task-relevant information while effectively filtering out noise. In fact, adding the conditioning on the task (U|T) is just slightly beneficial in terms of performance, with the exception of SVHN. The vast majority (on average around 90%) of alignment between task and residual comes from the head contributions, as witnessed by the similarity between the columns H and B. The optimized selection strategy (O) is extremely powerful, having the best score among them, and sometimes almost doubling the original model performances (B), showing how task-relevant information is already present in the residual, just hidden.
@@ 4.2 Spectral ResiDual Alignment
Given that: i) unit specialization is essentially encoded in the principal components (Section 3.2); ii) components of general enough data distributions (e.g., ImageNet) capture specialized behavior even on other datasets (Section 3.3); iii) retaining only task-aligned units is beneficial for image-text alignment (Section 4.1), we propose a method to directly filter information along the residual at the spectral level: ResiDual.
In short, we start from the decomposition formula for the residual stream (Equation 1) and allow anisotropic scaling of each unit representation. Specifically, given a unit representation ${\bm{X}}$, its corresponding principal component basis $\bm{\Phi}$, and associated mean $\bm{\mu}$, we define the ResiDual transformation of ${\bm{X}}$ as:
$\text{RD}_{\bm{\Phi},\bm{\mu}}({\bm{X}},\bm{\lambda})=\mathbf{\Phi}^{-1}\text{diag}(\bm{\lambda})\mathbf{\Phi}({\bm{X}}-\bm{\mu})^{T}\,,$ (6)
@@
${\bm{Y}}^{\prime}=\sum_{i=1}^{|{\mathbb{U}}|}\text{RD}_{\bm{\Phi_{i}},\bm{\mu_{i}}}({\bm{U}}_{i},\bm{\lambda}_{i})\,.$ (7)
In summary, ResiDual models a simple spectral anisotropic scaling of residual units that results in more complex dynamics in the output space. In this section, we extensively evaluate the effectiveness of this method across a variety of configurations, models and datasets.
@@
We evaluate ResiDual in 3 different configurations: i) RD, as presented in the original formulation; ii) RD, a version of ResiDual constrained on the number and type of units (we restrict it to heads) and the number of considered components (we truncate PCA bases to explain 90% of the variance). This is done because head units, and specifically their principal components, are strongly connected to specialization; iii) RDY, where the ResiDual transformation is applied directly on the output encoding ${\bm{Y}}$, to assess whether output components can already be aligned with the task. For all configurations, we select ImageNet as a reference for the PCA bases $\bm{\Phi}$ and unit means $\bm{\mu}$ that appear in the ResiDual Equation 6. Please refer to Table 8 for additional experiments with dataset-specific bases.
@@
For this experiment, we work with 3 CLIP-like models, BLIP-L, CLIP-L, and OpenCLIP-L (CLIP/OpenCLIP-B results can be found in the Appendix in Table 9 and Figure 21), and tune them on the 10 datasets employed in Section 4.1. All training runs use the Schedule-Free Adam optimizer (Defazio et al., 2024) with the automatic learning rate finder by Smith (2017), implemented in PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019). The maximum number of epochs is 30, with an early-stopping policy on the validation set accuracy with patience of 5 epochs.
@@
Result analysis
A comparative analysis of ResiDual against reference alignment methods is reported in Figure 5. We observe that a linear transformation of the output is sufficient to approximate full finetuning performance. The spectral residual transformation modelled by ResiDual attains comparable (and, in some cases, better) results than the linear aligner on all datasets. These statements hold true across all models. Specifically, the case of SVHN stands out: on this input dataset, ResiDual has an advantage of approximately 10% on all models. We hypothesize that this gap is due to the absence of task-relevant features at the output level. This is confirmed by the results in Table 2: on SVHN, RDY has the largest gap from RD, meaning that output components are not well aligned with the task. While having approximately 30% of the learnable parameters, RD achieves comparable results to RD, indicating that applying the ResiDual procedure to heads (and not MLPs) and their first principal components alone is enough. Moreover, we note that in the cases where the performance of the original model is already satisfactory (e.g., CIFAR-10), applying ResiDual does not compromise alignment.
@@ 5 Conclusions
ResiDual fundamentally works by extracting information from residual units already containing task-relevant principal components. We expect that when this assumption does not hold, ResiDual cannot recover the alignment, resulting in a significant drop in downstream performance. Moreover, our downstream tasks focused on zero-shot image classification. This implies considering only the alignment between [CLS] token and classifier, ignoring sequence-level information. However, preliminary results show high spectral compatibility between the [CLS] representations and the other patch tokens (Figure 12).
@@
The ResiDual formulation is based solely on the residual decomposition technique, which opens to its application across virtually any transformer architecture (with some preliminary results on unimodal encoders in Section A.3 ). Our findings show that ResiDual can be constrained at the coarse level to operate on a subset of units (i.e., attention heads) and at the fine level on specific principal components. Additional constraints to focus on the first few model layers create opportunities to enhance model inference, while the possibility to disable specific spectral components opens up fairness applications.
@@ Acknowledgements
Luca Bortolussi was supported by the PNRR project iNEST (Interconnected Nord-Est Innovation Ecosystem) funded by the European Union NextGenerationEU (Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS 00000043, CUP J43C22000320006).
@@ A.1 Residual decomposition
Transformers are residual networks made of stacked blocks that contain a Multi-Head Attention (MHA) layer and an MLP, both surrounded by a residual connection. All vision transformers we employ in this work abide by the “pre-norm” Wang et al. (2019) architecture, which slightly modified the original (Vaswani et al., 2017) by moving LayerNorm before MHA and MLP sub-layers (Figure 6). This modification implies that, at each layer, MHA and MLP sub-layers directly write their output representation to the residual stream. Hence, the final latent representation of the model is given by:
@@
In many transformer models, such as the ones employed in this work, the final residual encoding ${\bm{Y}}$ is not the final output of the model. For instance, in CLIP, ${\bm{Y}}$ is passed through a LayerNorm and then through a bias-less linear projection ${\bm{P}}$ that maps ViT encodings to the shared vision-language space: